OpenStudy (anonymous):

On segment AB, C is between A and B. Given AC = 12 and CB = 16, which is proportional to the ratio of (BC − AC) to AB?

OpenStudy (anonymous):

I would help you but I suck at geometry :/

OpenStudy (anonymous):

Are there options?

OpenStudy (anonymous):

it's either 1/8, 8, 7/1, or 1/7

OpenStudy (anonymous):

AB = 28... BC - AC = 4 so BC-AC : AB = 4/28 = 1/7

OpenStudy (anonymous):

Thank you guys!